# Foundation Models in Electrocardiogram: A Review

### v2

@@ 4.2 Pipeline
Predictive-based pre-training.
In predictive-based pre-training, the FM generates future tokens by conditioning each token on the sequence of previously generated tokens. When applied to ECG analysis, this method typically utilizes Transformer-based models to encode input ECG data, such as heartbeat sequences and image patches, into high-dimensional tokens. These sequential tokens with position information enable FMs to capture temporal and spatial dependencies, as well as long-range relationships across ECG embeddings, like arrhythmias spanning multiple timestamps of recording. Current research commonly employs LLMs like GPTs as pre-trained encoders to obtain robust feature representations for downstream tasks. For instance, zero-shot RAG [135] uses LLaMA2 and GPT3.5 to convert ECG signals into prompts and integrates expert domain knowledge to guide LLMs in diagnosing arrhythmia and sleep apnea. The specific capability of predictive-based pre-training methods on human instruction comprehension allows ECG-FMs to adapt to question-answering tasks, essentially acting as medical AI bots. SignalGPT [69], for example, offers an interactive interface for users to obtain essential information and fine-tunes ChatGPT by providing feedback. It uses ChatGPT as a comprehensive controller, determining the processing engine according to the signal types and analyzing the description of the given signal, based on pre-training on a massive ECG corpus.
Hybrid pre-training.
Several studies have effectively employed hybrid methods that combine contrastive-based, reconstruction-based, and predictive-based techniques in the pre-training process. Hybrid pre-training has been shown to significantly improve the efficiency and robustness of representation learning by integrating the advantages of these methods. First, hybrid pre-training can effectively mitigate the trade-off between sample diversity and the introduction of noise that generally accompanies data augmentation. For example, GCL [107] integrates denoising autoencoder and augmentation contrast, where the decoder reconstructs the augmented views of ECG samples to reduce the noise introduced by data augmentation. Furthermore, hybrid pre-training notably enhances the generalization capability of ECG-FMs by refining the multimodal representations of the samples. A compelling example is CQA-ESI [136], which adopts an LLM pipeline to generate detailed textual descriptions of ECGs and leverages contrastive and captioning loss to pre-train ECG encoders for enhanced representations. This approach demonstrates exceptional performance in zero-shot tasks.
3.2.2 Adaptation Techniques.
@@
Linear probe is the simplest and most commonly used strategy for adapting ECG-FMs to downstream tasks. This approach involves training a simple linear classifier, such as logistic regression (LR) or a fully-connected layer (FC), based on the feature representations learned by the pre-trained FM. During the adaptation process, only the weights of the linear classifier are updated through back-propagation, while the remaining FM parameters are kept frozen. The performance of this linear classifier can provide valuable insights into the usefulness of the learned representations for the task at hand. Consequently, it facilitates an assessment of the informational content within the model’s learned representations for a specific task or class, without the need for further fine-tuning of the model itself.
@@
N/A FC LP CAD PhysioNet/ CinC2021 PhysioNet/ CinC2021 ECG TS
@@
11footnotemark: 1Lead-agnostic SSL [90] wav2vec+CMSC ED FC LP CAC;
@@
22footnotemark: 2PIDA [55] MSDNN+MoCo N/A MSDNN F-FT CAD Cardiocloud W12L-ECG Cardiocloud W12L-ECG ECG TS
@@
33footnotemark: 3BIOT [130] Linear
@@
44footnotemark: 4WCR [80] wav2vec+CMSC EO FC LP LVEF;
@@
PC 66footnotemark: 6SSMs [82] S4 with FCE N/A MLP A-FT CAC CinC PTB-XL;
@@
M 77footnotemark: 7Adversarial masking [10] ResNet1d-18 N/A FC LP CAC;
@@
77footnotemark: 7Adversarial masking [11] ResNet1d-18 N/A FC LR CAC;
@@
88footnotemark: 8HeartBEiT [118] Dall-E+BEiT ED MLP F-FT LVEF;
@@
MTECG [143] Transformer ED FC A-FT AC Fuwai;
@@
99footnotemark: 9MassMIB [132] CNN+ViT EO FC F-FT CAD PTB-XL;
@@
Predictive AR Stacked Transformer [31] Transformer ED FC;
@@
Hybrid AE+AR 1010footnotemark: 10Auto-TTE [16] VQ-VAE DO N/A N/A Synthesis PTB-XL;
@@
  • a
    EO Encoder-only, DO Decoder-only, ED Encoder-decoder.
  • b
    AC Augmentation Contrast, MC Multimodal Contrast, AC Augmentation Contrast, AR Autoregressive-based, M Masking, AE Autoencoder-based, PC Prediction Contrast.
  • c
    FC Fully-connected Layer, LR Linear Regression, MLP Multilayer Perceptron.
  • d
    LP Linear Probe, F-FT Full Fine-tuning, P-FT Partial Fine-tuning, A-FT Additive Fine-tuning, ZS Zero-shot.
  • e
    CAC Cardiac Arrhythmia Classification, CAD Cardiac Abnormality Detection, MIC Myocardial Infarction Classification, Stress Stress Detection, BGL Blood Glucose Level Detection, Chatbot ECG Chatbot Assistant, Patient Patient Classification, Noise Noise Detection, ZSC Zero-shot Classification, Synthesis ECG Synthesis, Sleep stage Sleep Stage Classification, Sleep apnea Sleep Apnea Detection, OOD OOD Detection, Emotion Emotion Recognition, Blood glucose level Blood Glucose Level Monitoring, Denoising ECG Denoising, ACTN Abnormal Cardiac Troponin Diagnosis, LVSD LVSD Diagnosis, LVEF LVEF Diagnosis, HCM HCM Diagnosis, STEMI STEMI Diagnosis, SHDM Structural Heart Disease Murmurs Diagnosis, AF Atrial Fibrillation Diagnosis, LEF Low Ejection Fraction Diagnosis, Gender Gender Classification, Age Age Classification, BMI BMI Classification, Sex Sex Classification.
  • f
    TS Time Series, Plot ECG Plot, TF Time-Frequency Image, FD Frequency Domain.
  • 1
    https://github.com/Jwoo5/fairseq-signals
  • 2
    https://github.com/SMU-MedicalVision/ECG-Classfication
  • 3
    https://github.com/ycq091044/BIOT
  • 4
    https://github.com/bowang-lab/ECG-FM
  • 5
    https://github.com/Aiwiscal/CDCL-ECG
  • 6
    https://github.com/tmehari/ssm_ecg
  • 7
    https://github.com/jessica-bo/advmask_ecg
  • 8
    https://github.com/akhilvaid/HeartBEiT
  • 9
    https://github.com/ysxGitHub/MassMIB
  • 10
    https://github.com/TClife/text_to_ecg
5 Discussion
Despite the progress made in developing ECG-FMs, several critical challenges remain that hinder their widespread adoption and effective application in clinical settings:
Limited downstream tasks were tested. The predominant focus of ECG-FMs has been on classifying cardiac diseases, with the majority of studies (37 out of 43) concentrating on the detection of arrhythmias and general abnormalities. However, this has raised concerns about the thoroughness of model evaluation and the breadth of diseases being addressed. There remains a wealth of unexplored but common cardiac disease cohorts, including hyperkalemia [28], left ventricular systolic dysfunction [7], silent atrial fibrillation [91], and hypertrophic cardiomyopathy [53], which demand our attention in developing ECG-FM. Additionally, the heavy reliance on a limited number of ECG datasets for pre-training and validation presents a significant limitation. Current research heavily leans on public datasets like PTB-XL, Physionet, CPSC, Chapman, CODE, CinC, and MIMIC. However, the scarcity of certain cardiac conditions in these datasets leads to imbalanced categories, potentially introducing bias in the detection of rare abnormalities. Therefore, it is crucial to conduct additional downstream tasks using a more diverse set of ECG data to thoroughly explore the potential of FMs in the nuanced realm of ECG analysis.
@@ 5 Discussion
Model weights are not shared. Out of 46 studies on ECG-FM, only 3 papers [10, 130, 72] have shared their saved weight files of the pre-trained model. These papers were able to do so because they utilized public datasets. Notably, the LLM community has a strong tradition of supporting and sharing their pre-trained FMs, as evidenced by the LLaMA-series, Baichuan-series, and Qwen-series models. However, the ECG research community lacks a clear guide on how to share pre-trained weights due to data privacy and ethics concerns, which hampers the development of ECG-FMs. Additionally, data isolation remains a significant challenge, with hospitals being hesitant to share their limited ECG data and collaborate to build a comprehensive ECG database. As a result, there is a lack of diversity in available ECG data for researchers and cardiologists. Furthermore, while high-tech companies typically develop and share state-of-the-art FMs, hospitals and medical institutions struggle to make full use of their real-time ECG data due to limited computing resources. Therefore, it is crucial to encourage collaboration among these institutions to drive the development of ECG-FMs and fully leverage valuable ECG data.
@@
Data quality is not considered. LLMs typically employ stringent data quality filtering processes to ensure that they are trained on high-quality textual data. This approach enhances the models’ ability to generate accurate and coherent outputs by learning correct syntax, semantics, and contextual nuances from clean and reliable sources. However, when this methodology is directly applied to the development of ECG-FMs, it presents significant challenges. In real clinical practice, a substantial portion of ECG data is of suboptimal quality due to various factors such as patient movement, poor electrode placement, electrical interference, and physiological artifacts like muscle contractions or respiratory movements [39, 70, 4]. If ECG-FMs are trained exclusively on high-quality, artifact-free ECG recordings, they may lack the robustness to handle the noisy and diverse data commonly encountered in clinical environments. This limitation can lead to significant degradation in model performance when deployed in real-world settings, potentially resulting in inaccurate interpretations or missed detections of critical cardiac abnormalities.
@@
The larger the better? We don’t know. The question of whether larger models inherently lead to better performance in the realm of ECG-FMs remains unanswered. Empirical studies have undeniably shown that LLMs adhere to scaling laws—principles indicating that model performance improves predictably with increases in parameters, training data size, and computational resources [40]. These scaling laws suggest that as models grow, they can capture more intricate patterns and nuances within the data, leading to enhanced performance across various tasks. However, the applicability of these scaling laws to the ECG domain is still under investigation. ECG data possess unique characteristics, such as complex temporal dynamics and significant inter-individual variability, which may influence how model size affects performance. Unlike text data, ECG signals require models to interpret subtle physiological variations that might not benefit simply from increased parameter counts. Therefore, it is uncertain whether larger ECG models will provide better diagnostic accuracy or generalization capabilities.
ECG-FM may have the same problem as LLMs. LLMs often prove quite challenging to interpret, making it difficult to grasp the processes behind representation and output generation [140]. Despite various theoretical analyses attempting to understand pre-training and the generalization capabilities, the lack of a strong theoretical foundation remains a significant hurdle. As a result, LLMs may generate responses that are factually incorrect and misleading, yet sound plausible and coherent, a phenomenon referred to as LLM hallucinations [25]. This issue extends to ECG-FMs, posing a serious risk in the healthcare domain. In the sensitive field of ECG analysis, the potential for false knowledge or hallucinations can lead to misdiagnoses and subsequent symptoms. Moreover, ECG-FMs function as black boxes, generally providing limited insights into their decision-making processes, presenting a challenge for their practical application due to a lack of transparency and failure to meet regulatory compliance requirements. While researchers are working to address this issue, further research is necessary to enhance the fairness of ECG-FMs. Developing visualization tools that can spotlight relevant ECG features could be crucial in mitigating misinformation. Furthermore, incorporating explainable AI (XAI) techniques into ECG analysis will strengthen the trust between AI models and clinicians.
@@
When and how are ECG-FMs ready for cardiologists? While ECG-FMs have demonstrated impressive versatility and adaptability for various cardiac-related tasks, they are still in the experimental phase. Ongoing research is focused on enhancing their performance and interpretability for clinical use. Notably, existing ECG AI tools like CardioAI and IBM Watson [48] have already gained regulatory approvals and are being implemented into clinical practice with user-friendly interfaces. However, there is still progress to be made before ECG-FMs become widely accessible and seamlessly integrated into real-time workflows. On the other hand, how can ECG-FMs really help cardiologists? First, ECG-FMs can greatly aid in cardiology report generation, potentially improving the precision of reports. Furthermore, these advanced systems have the potential to support the detection of challenging or rare cardiac conditions, offering valuable insights and suggestions in cases where cardiologists may struggle to make a confident diagnosis.
6 Conclusions

### v3

@@ Introduction
Predictive-based pre-training
In predictive-based pre-training, the FM generates future tokens by conditioning each token on the sequence of previously generated tokens. When applied to ECG analysis, this method typically utilizes Transformer-based models to encode input ECG data, such as heartbeat sequences and image patches, into high-dimensional tokens. These sequential tokens with position information enable FMs to capture temporal and spatial dependencies, as well as long-range relationships across ECG embeddings, like arrhythmias spanning multiple timestamps of recording. Current research commonly employs LLMs like GPTs as pre-trained encoders to obtain robust feature representations for downstream tasks. For instance, zero-shot RAG [72] uses LLaMA2 and GPT-3.5 to convert ECG signals into prompts and integrates expert domain knowledge to guide LLMs in diagnosing arrhythmia and sleep apnea. The specific capability of predictive-based pre-training methods on human instruction comprehension allows ECG-FMs to adapt to question-answering tasks, essentially acting as medical AI bots. SignalGPT [69], for example, offers an interactive interface for users to obtain essential information and fine-tunes ChatGPT by providing feedback. It uses ChatGPT as a comprehensive controller, determining the processing engine according to the signal types and analyzing the description of the given signal, based on pre-training on a massive ECG corpus.
Hybrid pre-training
Several studies have effectively employed hybrid methods that combine contrastive-based, reconstruction-based, and predictive-based techniques in the pre-training process. Hybrid pre-training has been shown to significantly improve the efficiency and robustness of representation learning by integrating the advantages of these methods. First, hybrid pre-training can effectively mitigate the trade-off between sample diversity and the introduction of noise that generally accompanies data augmentation. For example, GCL [120] integrates denoising autoencoder and augmentation contrast, where the decoder reconstructs the augmented views of ECG samples to reduce the noise introduced by data augmentation. Furthermore, hybrid pre-training notably enhances the generalization capability of ECG-FMs by refining the multimodal representations of the samples. A compelling example is CQA-ESI [70], which adopts an LLM pipeline to generate detailed textual descriptions of ECG data and leverages contrastive and captioning loss to pre-train ECG encoders for enhanced representations. This approach demonstrates exceptional performance in zero-shot tasks.
Adaptation techniques
@@
Linear probe
Linear probe is one of the most commonly used strategies for adapting ECG-FMs to downstream tasks. This approach involves training a simple linear classifier, such as logistic regression (LR) or a fully-connected layer (FC), based on the feature representations learned by the pre-trained FM. During the adaptation process, only the weights of the linear classifier are updated through back-propagation, while the remaining FM parameters are kept frozen. The performance of this linear classifier can provide valuable insights into the usefulness of the learned representations for the task at hand. Consequently, it facilitates an assessment of the informational content within the model’s learned representations for a specific task or class, without the need for further fine-tuning of the model itself.
Fine-tuning
@@
N/A FC LP CAD PhysioNet;
CinC2021
PhysioNet;
CinC2021
ECG TS
11footnotemark: 1CLOCS [134] CNN1d N/A FC;
MLP
LP;
F-FT
CAC PhysioNet2020;
Chapman
PhysioNet2020;
Chapman;
PhysioNet2017;
Cardiology
ECG TS
CPT [16] ResNet1d N/A MLP F-FT mPAP;
Mortality
MGH MGH ECG;
Vital signs
TS
@@
22footnotemark: 2Lead-agnostic SSL [136] wav2vec+CMSC ED FC LP CAC;
@@
33footnotemark: 3PIDA [141] MSDNN+MoCo N/A MSDNN F-FT CAD Cardiocloud W12L-ECG Cardiocloud W12L-ECG ECG TS
@@
44footnotemark: 4BIOT [68] Linear
@@
55footnotemark: 5WCR [145] wav2vec+CMSC EO FC LP LVEF;
@@
JMCL [95] ResNet1d+
MedLlama3-JP-v2text
EO FC ZS ZSC UTH;
Mitsui
UTH;
Mitsui
ECG;
ECG reports
TS+Text
M3Bind [94] ECG-CLIP EO LoRA ZS ZSC MIMIC-CXR;
Quilt-1M
MESIDOR;
ODIR;
CheXpert;
RSNA;
MIMIC-ECG;
PTB-XL;
ICBEB;
SkinCancer;
Camelyon;
SICAPv2;
SkinTumor;
CT-RATE;
RAD-ChestCT
ECG;
Retina;
X-ray;
CT;
Pathology;
Reports
Image+Text
PC 77footnotemark: 7SSMs [126] S4 with FCE N/A MLP A-FT CAC CinC PTB-XL;
@@
M 88footnotemark: 8Advm [118] ResNet1d-18 N/A FC LP CAC;
@@
88footnotemark: 8Advm [119] ResNet1d-18 N/A FC LR CAC;
@@
99footnotemark: 9HeartBEiT [74] DALL-E+BEiT ED MLP F-FT LVEF;
@@
MTECG [76] Transformer ED FC A-FT CAC Fuwai;
@@
ECGBERT [90] BERT EO MLP F-FT AF;
Sleep apnea;
Heartbeat;
Patient
MIMIC-III;
PTB-XL;
Georgia;
CPSC-2018
MIT-BIH ECG TS
MultiMAE [148] MAE N/A FC LP Sleep;
Age;
Arousal
PhysioNet2018 PhysioNet2018 ECG;
EEG;
EOG;
EMG
TS
1010footnotemark: 10MassMIB [128] CNN+ViT EO FC F-FT CAD PTB-XL;
@@
Predictive AR SK-TF [75] Transformer ED FC;
@@
CDE-MedED [108] ChatGPT;
DALL-E
DO N/A ZS Synthesis N/A N/A ECG;
X-ray
Image
DECG [107] DALL-E DO N/A ZS Synthesis N/A N/A ECG ECG plot
CBPM-LLaMA [103] LLaMA3 DO N/A F-FT CBPM N/A CAS-BP ECG;
PPG
TS
ECG-LM [99] BioMedGPT-LM;
GPT-3.5
DO FC F-FT ECG QA;
CAD;
ZSC
N/A PTB-XL;
PTB-XL+;
ECG-QA
ECG TS+Text
1111footnotemark: 11ECG-LLM [102] LLaMA2;
LLaMA3
DO MLP F-FT Forecasting;
Imputation
N/A PhysioNet2020 ECG TS
1212footnotemark: 12ECG-Expert-QA [98] GPT-4o;
DeepSeek-v3;
Qwen-2.5;
MiniMind-2
DO N/A ZS ECG QA N/A ECG-Expert-QA ECG;
ECG reports
TS+Text
Hybrid AE+AR 1313footnotemark: 13Auto-TTE [20] VQ-VAE DO N/A N/A Synthesis PTB-XL;
@@
AR+MC 1414footnotemark: 14GEM [112] ECG-CoCa;
GPT-4o;
SFT-LLaVA;
SFT-PULSE
ED MLP LoRA Grounded
understanding
MIMIC-IV-ECG;
PULSE;
ECG-Bench
MIMIC-IV-ECG;
PULSE;
ECG-Bench
ECG TS+Plot+Text
1515footnotemark: 15ECG-Chat [111] ECG-CoCa;
GPT-4o
ED FC LoRA Report
generation
N/A ECG-Instruct ECG;
ECG reports
TS+Text
Discussion
Our discussion highlights both challenges and opportunities. In addition to addressing the key barriers hindering the clinical adoption of ECG-FMs, we underscore important findings that advance ECG interpretation and support broader healthcare tasks.
Multimodal inputs help understand ECG data.
Recent advances in ECG-FMs demonstrate that performance depends not only on model architecture but also on input representation and the richness of available information. While many models directly process raw ECG signals, others exploit transformed representations such as ECG plots, spectrograms, or time–frequency images. These variations underscore the need to carefully consider data format, as it shapes both model design and performance. For example, MTECG [76] converts ECG segments into image-like patches and applies masked autoencoders to reduce redundancy and enhance wave-shape feature learning, whereas BIOT [68] employs a frequency-domain tokenization strategy using Fourier-based energy vectors to capture spectral dynamics and enable robust cross-dataset pre-training. However, whether raw signals or transformed representations provide a superior foundation for ECG representation learning remains an open question.
Beyond input representation, multimodal integration has emerged as a promising approach to enrich ECG understanding. In addition to ECG, several studies recent studies incorporate related physiological signals (e.g., PPG, EMG) and complementary information (e.g., EHRs, spectrograms). By combining multiple modalities, these frameworks capture a more holistic view of cardiovascular function and patient state. Empirical evidence shows that multimodal models consistently outperform unimodal counterparts across downstream tasks and model architectures. For instance, SSL-KD [121] fuses ECG series with spectrograms for arrhythmia classification, boosting accuracy from 48.8% to 48.9% and F1-score from 61.4% to 62.1% on PhysioNet2020; ECG-PCG-MAE [129] achieves an average Area Under the ROC Curve (AUC) gain of 5.5% for LEF detection compared to ECG-only or PCG-only models; GEM [112] integrates raw ECG, transformed images, and textual instructions to yield a 5.8% improvement for grounded abnormality detection on PTB-XL and CODE. Overall, these findings suggest that multimodal frameworks not only enhance predictive performance but also strengthen interpretability and alignment with clinical reasoning. Each modality provides unique and complementary cues: for example, spectrograms highlight frequency-domain patterns such as rhythm regularity or noise artifacts that may be overlooked in raw ECG; PCG captures mechanical heart sounds indicative of valve function; PPG reflects peripheral vascular responses and oxygenation dynamics; and EHRs encode semantic knowledge of patient history, comorbidities, and treatments. By unifying these heterogeneous sources, ECG-FMs can learn richer latent representations that jointly capture electrophysiological, hemodynamic, and contextual dimensions of cardiovascular health. This integration not only mitigates sensitivity to noise and artifacts in any single modality, but also aligns model reasoning more closely with real-world clinical practice, where clinicians routinely synthesize heterogeneous evidence from multiple streams to reach informed decisions.
ECG-FMs still face barriers to clinical workflow integration.
Despite the strong performance across benchmark tasks, most ECG-FMs are still evaluated in controlled, offline settings and have not yet been embedded into routine clinical workflows. However, real-world deployment requires more than predictive accuracy. Models should integrate with with EHRs and hospital information systems, provide interpretable outputs consistent with clinical reasoning, and function reliably under real-time and noisy conditions [149, 150]. These challenges, along with regulatory compliance, data privacy, and usability concerns, remain underexplored, leaving a translational gap between research prototypes and clinical adoption. Nevertheless, ECG-FMs hold considerable promise for integration into practice. For example, they could assist clinicians by generating preliminary diagnostic reports, prioritizing high-risk cases for review, or combining ECG-derived representations with multimodal patient data to inform treatment decisions [151, 152]. Therefore, future research should emphasize prospective validation in clinical environments, development of user-centered interfaces, and collaboration with regulatory bodies to ensure safe and effective deployment. Addressing these challenges will be essential for ECG-FMs to evolve from promising research tools into practical decision-support systems capable of enhancing patient care.
Limited downstream tasks are tested.
As illustrated in Figure 5, we categorize the applications of ECG-FMs into 11 types: (1) general cardiac arrhythmia classification/detection (CAC/CAD), (2) stress monitoring, (3) sleep monitoring (e.g., sleep stage and sleep apnea), (4) special cardiac diseases (e.g., LVSD, LVEF, ACTN, HCM), (5) patient information identification (e.g., age and gender), (6) zero-shot inference, (7) ECG synthesis, (8) blood parameter estimation (e.g., mPAP and BGL), (9) ECG question answering, (10) signal refinement (e.g., imputation and denoising), and (11) other tasks (e.g., emotion and arousal classification). Our analysis reveals that the predominant focus has been on classification tasks, with over 40% of studies addressing CAC or CAD, while nearly 20% target patient information identification and zero-shot classification. In contrast, fewer than 10% of studies apply ECG-FMs to other cardiac diseases, raising concerns about the comprehensiveness of model evaluation and the breadth of clinical conditions being addressed. Notably, there remain many common but underexplored cardiac disease cohorts, such as hyperkalemia [153] and hypertrophic cardiomyopathy [154], which warrant greater attention in ECG-FM development. Another limitation lies in the heavy reliance on a small set of public ECG datasets, such as PTB-XL, Chapman, CODE, CinC, and MIMIC. The scarcity of certain conditions in these datasets results in imbalanced categories, potentially biasing the detection of rare abnormalities. To address these limitations, future research should pursue a broader set of downstream tasks and leverage more diverse ECG datasets to fully unlock the potential of foundation models in advancing nuanced and clinically meaningful ECG analysis.
Figure 5: Types of downstream tasks in the reviewed papers.
Model weights are not shared.
@@ Discussion
Out of 60 studies on ECG-FM, only 3 papers [68, 116, 145] have shared their saved weight files of the pre-trained model. These papers were able to do so because they utilized public datasets. Notably, the LLM community has a strong tradition of supporting and sharing their pre-trained FMs, as evidenced by the LLaMA-series, Baichuan-series, and Qwen-series models. However, the ECG research community lacks a clear guide on how to share pre-trained weights due to data privacy and ethics concerns, which hampers the development of ECG-FMs. Additionally, data isolation remains a significant challenge, with hospitals being hesitant to share their limited ECG data and collaborate to build a comprehensive ECG database. As a result, there is a lack of diversity in available ECG data for researchers and cardiologists. Furthermore, while high-tech companies typically develop and share state-of-the-art FMs, hospitals and medical institutions struggle to make full use of their real-time ECG data due to limited computing resources. Therefore, it is crucial to encourage collaboration among these institutions to drive the development of ECG-FMs and fully leverage valuable ECG data.
Data quality is not considered.
@@
LLMs typically employ stringent data quality filtering processes to ensure that they are trained on high-quality textual data. This approach enhances the models’ ability to generate accurate and coherent outputs by learning correct syntax, semantics, and contextual nuances from clean and reliable sources. However, when this methodology is directly applied to the development of ECG-FMs, it presents significant challenges. In real clinical practice, a substantial portion of ECG data is of suboptimal quality due to various factors such as patient movement, poor electrode placement, electrical interference, and physiological artifacts like muscle contractions or respiratory movements [155, 156, 157]. If ECG-FMs are trained exclusively on high-quality, artifact-free ECG recordings, they may lack the robustness to handle the noisy and diverse data commonly encountered in clinical environments. This limitation can lead to significant degradation in model performance when deployed in real-world settings, potentially resulting in inaccurate interpretations or missed detections of critical cardiac abnormalities.
Benchmarking framework remains unstandardized.
A critical limitation in the current development of ECG-FMs lies in the lack of standardized benchmarking frameworks. Although recent studies have demonstrated strong performance across diverse datasets and tasks, results are typically reported using heterogeneous datasets, evaluation metrics, and experimental protocols, which hinders fair comparison across models and complicates the assessment of generalizability [158]. This fragmentation not only obscures the relative strengths and weaknesses of different approaches but also raises concerns regarding robustness when models are applied to heterogeneous patient populations and real-world clinical settings. Establishing standardized evaluation suites, incorporating cross-dataset testing protocols, unified performance metrics, and task-specific benchmarks, would thus provide a more rigorous foundation for structuring future ECG-FM research. While no comprehensive benchmarking standard currently exists for ECG-FMs, collaborative efforts between the clinical and machine learning communities could play a pivotal role in curating representative datasets and defining shared evaluation criteria.
The larger the better? We don’t know.
@@
The question of whether larger models inherently lead to better performance in the realm of ECG-FMs remains unanswered. Empirical studies have undeniably shown that LLMs adhere to scaling laws—principles indicating that model performance improves predictably with increases in parameters, training data size, and computational resources [159]. These scaling laws suggest that as models grow, they can capture more intricate patterns and nuances within the data, leading to enhanced performance across various tasks. However, the applicability of these scaling laws to the ECG domain is still under investigation. ECG data possess unique characteristics, such as complex temporal dynamics and significant inter-individual variability, which may influence how model size affects performance. Unlike text data, ECG signals require models to interpret subtle physiological variations that might not benefit simply from increased parameter counts. Therefore, it is uncertain whether larger ECG models will provide better diagnostic accuracy or generalization capabilities.
ECG-FMs may have the same problem as LLMs.
LLMs are often difficult to interpret, which complicates efforts to elucidate the mechanisms underlying their representation learning and output generation [160]. Despite ongoing theoretical efforts, a comprehensive understanding of their pre-training dynamics and generalization abilities remains limited. AS a result, LLMs are prone to producing factually incorrect yet coherent and plausible outputs, a phenomenon commonly known as hallucination [161]. Similar concerns arise for ECG-FMs, where the generation of misleading information poses serious risks in clinical contexts. In ECG analysis, such hallucinations may manifest as unreliable diagnostic cues, potentially leading to misinterpretation and adverse patient outcomes. This problem is compounded by the black-box nature of ECG-FMs, which provide limited transparency into their decision-making processes and hinder both clinical adoption and regulatory approval. To address these limitations, explainable AI (XAI) techniques have emerged as promising solutions. For example, Grad-CAM [162] can highlight salient temporal or spatial features in ECG signals, while SHAP [163] can attribute predictive importance to specific waveform segments. By offering visual or quantitative explanations aligned with clinical reasoning, these methods improve interpretability and foster trust. Although their use in ECG analysis remains at an early stage, integrating XAI frameworks into ECG-FMs could play a pivotal role in mitigating hallucination risks, enhancing transparency, and enabling safer deployment in real-world healthcare environments.
When and how are ECG-FMs ready for cardiologists?
@@
While ECG-FMs have demonstrated impressive versatility and adaptability for various cardiac-related tasks, they are still in the experimental phase. Ongoing research is focused on enhancing their performance and interpretability for clinical use. Notably, existing ECG AI tools like CardioAI and IBM Watson [164] have already gained regulatory approvals and are being implemented into clinical practice with user-friendly interfaces. However, there is still progress to be made before ECG-FMs become widely accessible and seamlessly integrated into real-time workflows. On the other hand, how can ECG-FMs really help cardiologists? First, ECG-FMs can greatly aid in cardiology report generation, potentially improving the precision of reports. Furthermore, these advanced systems have the potential to support the detection of challenging or rare cardiac conditions, offering valuable insights and suggestions in cases where cardiologists may struggle to make a confident diagnosis.
Conclusions